Would you please solve some equations for me for a discussion questions.
1. 5 y -1 >10
2. 6-3x>6
3. 4t >4 + 3 t
I do not get any answers when I submit questions, please this is important and I need help in alegebra. I am having some problems understanding. Thank you very much I appreciate any help you could help me with. Please show work as well it helps get points as well.

Answer by lwsshak3(11628)

1. 5 y -1 >10
5y>10+1=11
y>11/5
solution:
(11/5,∞)
..
2. 6-3x>6
-3x>6-6=0
-x>0
multiply both sides by (-1) and reverse inequality sign
x<0
solution
(-∞,0)
..
3. 4t >4 + 3 t
4t-3t>4
t>4
solution
(4,∞)